OpenStudy (selena2345):

What is the difference 6 3/12- 2 2/3?

OpenStudy (anonymous):

so the first one = 75/12 and the second is 8/3. you find the common denominator, which is twelve. so you multiply the top and bottom of the second one by 4. so that = 32/12. and then 75/12-32/12 = 43/12

OpenStudy (anonymous):

6 3/12- 2 2/3 = 75/12 − 8/3 = ((75 × 3) − (8 × 12)) / (12 × 3) = (225 - 96) / 36 = 129/36 = 43/12 Simplified to: 3 7/12